# MS-0001-1922.65

### v2

@@ 5.3 Impact of Self-Preference Bias in Algorithmic Hiring
Taken together, these findings show that self-preference bias is not simply a statistical artifact but a systematic force that can distort hiring outcomes, narrow candidate diversity, and undermine the fairness and reliability of AI-mediated recruitment. These risks highlight the need for safeguards and mitigation strategies before LLM-based hiring systems are deployed at scale.
@@ 6.1 Mechanism: Self-Recognition
Self-Preference Bias (%) 88 84 78
@@
Self-Preference Bias (%) 48 24 58
Absolute Decrease (pp) $\downarrow$ 40 60 20
Relative Decrease (%) $\downarrow$ 45.4 71.4 25.6
@@
Self-Preference Bias (%) 32 26 34
Absolute Decrease (pp) $\downarrow$ 56 58 44
Relative Decrease (%) $\downarrow$ 63.6 69.0 56.4
@@ 6.2 System Prompting
Applying this prompting strategy across evaluator models leads to consistent reductions in self-preference bias. For example, GPT-4o’s LLM-vs-Human bias decreases from $92\%$ to $48\%$, and DeepSeek-V3’s from $78\%$ to $58\%$, after controlling for resume quality (Table 5). These results indicate that the bias is not hardwired into model architecture but can be shaped by context and instruction, providing evidence that explicit prompts can partially disrupt self-recognition.
@@ 6.3 Majority Voting Ensemble
This mitigation strategy proves highly effective. Across all three LLM models tested, the average LLM-vs-Human comparisons self-preference bias can drop by over $50\%$. For example, as it is shown in Table 5, GPT-4o’s bias is reduced from $88\%$ to $32\%$, and LLaMA-3.3-70B’s from $84\%$ to $26\%$. These results demonstrate that both system prompting and majority voting offer robust and scalable approaches to mitigation, particularly in high-stakes applications where fairness is critical.

### v3

@@ 6.1 Mechanism: Self-Recognition
Self-Preference Bias (%) 82 79 72
@@
Self-Preference Bias (%) 61 30 60
Absolute Decrease (pp) $\downarrow$ 21 49 12
Relative Decrease (%) $\downarrow$ 26 62 17
@@
Self-Preference Bias (%) 30 23 29
Absolute Decrease (pp) $\downarrow$ 52 56 43
Relative Decrease (%) $\downarrow$ 63 71 60
@@ 6.2 System Prompting
Applying this prompting strategy across evaluator models leads to consistent reductions in self-preference bias. For example, GPT-4o’s LLM-vs-Human bias decreases from $82\%$ to $61\%$, and LLaMA-3.3-70B’s from $79\%$ to $30\%$, after controlling for resume quality (Table 2). These results indicate that the bias is not hardwired into model architecture but can be shaped by context and instruction, providing evidence that explicit prompts can partially disrupt self-recognition.
@@ 6.3 Majority Voting Ensemble
This mitigation strategy proves highly effective. Across all three LLM models tested, the average LLM-vs-Human comparisons self-preference bias can drop by over $50\%$. For example, as it is shown in Table 2, GPT-4o’s bias is reduced from $82\%$ to $30\%$, and LLaMA-3.3-70B’s from $79\%$ to $23\%$. These results demonstrate that both system prompting and majority voting offer robust and scalable approaches to mitigation, particularly in high-stakes applications where fairness is critical.